OpenStudy (anonymous):

The ordered pair (2, 0.6) is from an inverse variation. Find the constant of variation. a 0.3 B. 3.333 C. 1.2 D. 2.6

jimthompson5910 (jim_thompson5910):

wait, is this really a true/false question?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

but it says to find the constant of variation

OpenStudy (anonymous):

oh sorry i made a mistake. the options were a 0.3 B. 3.333 C. 1.2 D. 2.6

jimthompson5910 (jim_thompson5910):

solve for k y = k/x 0.6 = k/2

OpenStudy (anonymous):

1.2?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks again!

jimthompson5910 (jim_thompson5910):

np